OpenStudy (anonymous):

What would y be? ^3√(3y+3)=3

OpenStudy (anonymous):

Well the idea is to isolate y right? What would be your first step?

OpenStudy (anonymous):

You have a cube root. How would you get rid of that?

OpenStudy (anonymous):

For instance, if you have a square root you have to square to get rid of the square root.

OpenStudy (anonymous):

@francescazares2012 : You there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So I would cube both sides?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Is my answer 2?

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

It didn't work. :/

OpenStudy (anonymous):

the answer is 8! :D

OpenStudy (anonymous):

Ohh wait. I am dumb.

OpenStudy (anonymous):

3^3 is 27 >.< .

OpenStudy (anonymous):

thanks for your help! I was having so much trouble

OpenStudy (anonymous):

Yeah it's 8 :P .

OpenStudy (anonymous):

could you help me with another?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

√(x+10)=x+10

OpenStudy (anonymous):

Okay, square both sides again :) .

OpenStudy (anonymous):

okay so now i have x+10=(x+10)^2

OpenStudy (anonymous):

Can you expand the (x+10)^2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):